# Clearer Frames, Anytime: Resolving Velocity Ambiguity in Video Frame Interpolation

### v2

@@ Abstract
Existing video frame interpolation (VFI) methods blindly predict where each object is at a specific timestep $t$ (“time indexing”), which struggles to predict precise object movements. Given two images of a baseball, there are infinitely many possible trajectories: accelerating or decelerating, straight or curved. This often results in blurry frames as the method averages out these possibilities. Instead of forcing the network to learn this complicated time-to-location mapping implicitly together with predicting the frames, we provide the network with an explicit hint on how far the object has traveled between start and end frames, a novel approach termed “distance indexing”. This method offers a clearer learning goal for models, reducing the uncertainty tied to object speeds. We further observed that, even with this extra guidance, objects can still be blurry especially when they are equally far from both input frames (i.e., halfway in-between), due to the directional ambiguity in long-range motion. To solve this, we propose an iterative reference-based estimation strategy that breaks down a long-range prediction into several short-range steps. When integrating our plug-and-play strategies into state-of-the-art learning-based models, they exhibit markedly sharper outputs and superior perceptual quality in arbitrary time interpolations, using a uniform distance indexing map in the same format as time indexing. Additionally, distance indexing can be specified pixel-wise, which enables temporal manipulation of each object independently, offering a novel tool for video editing tasks like re-timing. The code is available at https://zzh-tech.github.io/InterpAny-Clearer/.
Keywords:
Frame interpolation Manipulation Velocity ambiguity
1 Introduction
@@
Yet, in both cases, the unsampled blank between the two frames, such as the motion between a ball’s starting and ending points, presents infinite possibilities. The velocities of individual objects within these frames remain undefined, introducing a velocity ambiguity, a myriad of plausible time-to-location mappings during training. We observed that velocity ambiguity is a primary obstacle hindering the advancement of learning-based VFI: Models trained using aforementioned time indexing receive identical inputs with differing supervision signals during training. As a result, they tend to produce blurred and imprecise interpolations, as they average out the potential outcomes.
@@
$\boxed{I_{t}=\mathcal{F}\left(I_{0},I_{1},\text{explicit motion hint}\right)\;\;\;\Rightarrow\;\;\;I_{t}=\mathcal{F}\left(I_{0},I_{1},D_{t}\right).}$ (2)
@@
In summary, our key contributions are: 1) Proposing distance indexing and iterative reference-based estimation to address the velocity ambiguity and enhance the capabilities of arbitrary time interpolation models; 2) Conducting comprehensive validation of the efficacy of our plug-and-play strategies across a range of state-of-the-art learning-based models. 3) Presenting an unprecedented manipulation method that allows for customized interpolation of any object.
@@
2 Related Work
2.1 Video frame interpolation
General overview. Numerous VFI solutions rely on optical flows to predict latent frames. Typically, these methods warp input frames forward or backward using flow calculated by off-the-shelf networks like  [37, 6, 12, 38] or self-contained flow estimators like [11, 50, 21]. Networks then refine the warped frame to improve visual quality. SuperSlomo [14] uses a linear combination of bi-directional flows for intermediate flow estimation and backward warping. DAIN [1] introduces a depth-aware flow projection layer for advanced intermediate flow estimation. AdaCoF [19] estimates kernel weights and offset vectors for each target pixel, while BMBC [29] and ABME [30] refine optical flow estimation. Large motion interpolation is addressed by XVFI [34] through a recursive multi-scale structure. VFIFormer [24] employs Transformers to model long-range pixel correlations. IFRNet [18], RIFE [11], and UPR-Net [15] employ efficient pyramid network designs for high-quality, real-time interpolation, with IFRNet and RIFE using leakage distillation losses for flow estimation. Recently, more advanced network modules and operations are proposed to push the upper limit of VFI performance, such as the transformer-based bilateral motion estimator of BiFormer [28], a unifying operation of EMA-VFI [50] to explicitly disentangle motion and appearance information, and bi-directional correlation volumes for all pairs of pixels of AMT [21]. On the other hand, SoftSplat [26] and M2M [10] actively explore the forward warping operation for VFI.
@@ 2.1 Video frame interpolation
Learning paradigms. One major thread of VFI methods train networks on triplet of frames, always predicting the central frame. Iterative estimation is used for interpolation ratios higher than $\times$2. This fixed-time method often accumulates errors and struggles with interpolating at arbitrary continuous timesteps. Hence, models like SuperSloMo [14], DAIN [1], BMBC [29], EDSC [4], RIFE [11], IFRNet [18], EMA-VFI [50], and AMT [21] have adopted an arbitrary time interpolation paradigm. While theoretically superior, the arbitrary approach faces challenges of more complicated time-to-position mappings due to the velocity ambiguity, resulting in blurred results. This study addresses velocity ambiguity in arbitrary time interpolation and offers solutions.
Prior work by Zhou et al. [55] identified motion ambiguity and proposed a texture consistency loss to implicitly ensure interpolated content resemblance to given frames. In contrast, we explicitly address velocity ambiguity and propose solutions. These innovations not only enhance the performance of arbitrary time VFI models but also offer advanced manipulation capabilities.
Segment anything The emergence of Segment Anything Model (SAM) [17] has marked a significant advancement in the realm of zero-shot segmentation, enabling numerous downstream applications including video tracking and segmentation [48], breakthrough mask-free inpainting techniques [49], and interactive image description generation [40]. By specifying the distance indexing individually for each segment, this work introduces a pioneering application to this growing collection: Manipulated Interpolation of Anything.
3 Velocity Ambiguity
@@
So far, we have been discussing the ambiguity for the fixed time interpolation paradigm, in which $t$ is set by default to $0.5$. For arbitrary time interpolation, the ambiguity becomes more pronounced: Instead of predicting a single timestep, the network is expected to predict a continuum of timesteps between 0 and 1, each having a multitude of possibilities. This further complicates learning. Moreover, this ambiguity is sometimes referred to as mode averaging, which has been studied in other domains [41]. See supplemental materials for details.
4 Disambiguation Strategies
@@ 4.1 Distance indexing
During inference, the algorithm does not have access to the exact distance map since $I_{t}$ is unknown. In practice, we notice it is usually sufficient to provide a uniform map $D_{t}=t$, similar to time indexing. Physically this encourages the model to move each object at constant speeds along their trajectories. We observe that constant speed between frames is a valid approximation for many real-world situations. In Sec. 5, we show that even though this results in pixel-level misalignment with the ground-truth frames, it achieves significantly higher perceptual scores and is strongly preferred in the user study. Precise distance maps can be computed from multiple frames, which can potentially further boost the performance. See a detailed discussion in supplementary materials.
4.2 Iterative reference-based estimation
@@
Importantly, in every iteration, we consistently use the starting and ending frames as reliable appearance references, preventing divergence of uncertainty in each step. By dividing a long step into shorter steps, the uncertainty in each step is reduced, as shown in Fig. 3 (b). While fixed time models also employ an iterative method in a bisectioning way during inference, our strategy progresses from near to far, ensuring more deterministic trajectory interpolation. This reduces errors and uncertainties tied to a single, long-range prediction. See more on the rationale for solving ambiguities in supplemental materials.
@@
Figure 5: Comparison of qualitative results. [T]: original arbitrary time VFI models using time indexing. [D]: models using our distance indexing. [T,R]: models using time indexing with iterative reference-based estimation. [D,R]: models using both strategies. All models use uniform maps. Zoom in for a closer look.
5 Experiment
5.1 Implementation
@@
We leveraged the plug-and-play nature of our distance indexing and iterative reference-based estimation strategies to seamlessly integrate them into influential arbitrary time VFI models such as RIFE [11] and IFRNet [18], and state-of-the-art models including AMT [21] and EMA-VFI [50]. We adhere to the original hyperparameters for each model for a fair comparison and implement them with PyTorch [31]. For training, we use the septuplet dataset from Vimeo90K [47]. The septuplet dataset comprises 91,701 seven-frame sequences at $448\times 256$, extracted from 39,000 video clips. For evaluation, we use both pixel-centric metrics like PSNR and SSIM [42], and perceptual metrics such as reference-based LPIPS [51] and non-reference NIQE [25]. Concerning the iterative reference-based estimation strategy, $D_{ref}$ during training is calculated from the optical flow derived from ground-truth data at a time point corresponding to a randomly selected reference frame, like $t/2$. In the inference phase, we similarly employ a uniform map for reference, for example, setting $D_{ref}=t/2$. See our source code for details.
5.2 Qualitative comparison
@@
5.3 Quantitative comparison
Convergence curves.
To further substantiate the efficacy of our proposed strategies, we also conducted a quantitative analysis. Fig. 6 shows the convergence curves for different model variants, i.e., $[T]$, $[D]$, and $[D,R]$. The observed trends are consistent with our theoretical analysis from Sec. 4, supporting the premise that by addressing velocity ambiguity, both distance indexing and iterative reference-based estimation can enhance convergence limits.
Comparison on Vimeo90K septuplet dataset.
In Tab. 1, we provide a performance breakdown for each model variant. The models $[D]$ and $[D,R]$ in the upper half utilize ground-truth distance guidance, which is not available at inference in practice. The goal here is just to show the achievable upper-bound performance. On both pixel-centric metrics such as PSNR and SSIM, and perceptual measures like LPIPS and NIQE, the improved versions $[D]$ and $[D,R]$ outperform the base model $[T]$. Notably, the combined model $[D,R]$ using both distance indexing and iterative reference-based estimation strategies performs superior in perceptual metrics, particularly NIQE. The superior pixel-centric scores of model $[D]$ compared to model $[D,R]$ can be attributed to the indirect estimation (2 iterations) in the latter, causing slight misalignment with the ground-truth, albeit with enhanced details.
Table 1: Comparison on Vimeo90K septuplet dataset. $[T]$ denotes the method trained with traditional arbitrary time indexing paradigm. $[D]$ and $[R]$ denote the distance indexing paradigm and iterative reference-based estimation strategy, respectively. $[R]$ uses 2 iterations by default. $[\cdot]_{u}$ denotes inference with uniform map as time indexes. We utilize the first and last frames as inputs to predict the rest five frames. The bold font denotes the best performance in cases where comparison is possible. While the gray font indicates that the scores for pixel-centric metrics, PSNR and SSIM, are not calculated using strictly aligned ground-truth and predicted frames.

### v3

# Disambiguation for Video Frame Interpolation

@@ Abstract
Existing video frame interpolation (VFI) methods blindly predict where each object is at a specific timestep $t$ (“time indexing”), which struggles to predict precise object movements. Given two images of a baseball, there are infinitely many possible trajectories: accelerating or decelerating, straight or curved. This often results in blurry frames as the method averages out these possibilities. Instead of forcing the network to learn this complicated time-to-location mapping implicitly together with predicting the frames, we provide the network with an explicit hint on how far the object has traveled between start and end frames, a novel approach termed “distance indexing”. This method offers a clearer learning goal for models, reducing the uncertainty tied to object speeds. We further observed that, even with this extra guidance, objects can still be blurry especially when they are equally far from both input frames (i.e., halfway in-between), due to the directional ambiguity in long-range motion. To solve this, we propose an iterative reference-based estimation strategy that breaks down a long-range prediction into several short-range steps. When integrating our plug-and-play strategies into state-of-the-art learning-based models, they exhibit markedly sharper outputs and superior perceptual quality in arbitrary time interpolations, using a uniform distance indexing map in the same format as time indexing without requiring extra computation. Furthermore, we demonstrate that if additional latency is acceptable, a continuous map estimator can be employed to compute a pixel-wise dense distance indexing using multiple nearby frames. Combined with efficient multi-frame refinement, this extension can further disambiguate complex motion, thus enhancing performance both qualitatively and quantitatively. Additionally, the ability to manually specify distance indexing allows for independent temporal manipulation of each object, providing a novel tool for video editing tasks such as re-timing. The code is available at https://zzh-tech.github.io/InterpAny-Clearer/.
Index Terms:
Video frame interpolation, Temporal super-resolution, Disambiguation, Video editing
I Introduction
@@
Yet, in both cases, the unsampled blank between the two frames, such as the motion between a ball’s starting and ending points, presents infinite possibilities. The velocities of individual objects within these frames remain undefined, introducing a velocity ambiguity, a myriad of plausible time-to-location mappings during training. Incorporating additional neighboring frames as input [8] can partially restrict the solution space by imposing constraints on motion trajectories, but it cannot fully resolve the ambiguity. We observed that velocity ambiguity is a primary obstacle hindering the advancement of learning-based VFI: Models trained using aforementioned time indexing receive identical inputs with differing supervision signals during training. As a result, they tend to produce blurred and imprecise interpolations, as they average out the potential outcomes.
@@
$\boxed{I_{t}=\mathcal{F}\left(I_{0},I_{1},\text{motion hint}\right)\;\Rightarrow\;I_{t}=\mathcal{F}\left(I_{0},I_{1},D_{t}\right).}$ (2)
@@
When using more than two input frames [8], nearby frames offer additional constraints that facilitate the computation of a pixel-wise dense distance map. The methods discussed so far employ a uniform distance map because a deterministic distance map cannot be derived from only two frames. Inspired by continuous parametric optical flow estimation [12], we utilize cubic B-splines and neural ordinary differential equations to estimate a dense distance map from four input frames, which improves our model’s performance across both perceptual and pixel-centric metrics. Furthermore, we make a trainable copy of the original interpolator architecture to refine the initial two-frame interpolation results using information from two additional frames $I_{-1}$ and $I_{2}$. This multi-frame refiner module is intended to fully harness the potential of additional frames, thereby enhancing multi-frame interpolation quality.
@@
In summary, our key contributions are: 1) Proposing distance indexing and iterative reference-based estimation to address the velocity ambiguity and enhance the capabilities of arbitrary time interpolation models; 2) Presenting an unprecedented manipulation method that allows for customized interpolation of any object. 3) Adopting a continuous distance map estimator and proposing a multi-frame fusion architecture to enhance interpolation quality across both perceptual and pixel-centric metrics.
A preliminary version of this work was presented in [13], where we focus on using a uniform distance map during inference because an accurate, pixel-wise distance map cannot be reliably calculated from two frames. Although using a uniform distance map produced plausible results with better perceptual quality, the predictions did not align with the ground truth, resulting in lower performance on metrics such as PSNR and SSIM. In this paper, we address this limitation by using multiple frames (more than two) as input and introduce a continuous distance map estimator that approximates the map from nearby frames. We also present a simple yet effective multi-frame refiner architecture for video frame interpolation. Extensive experiments demonstrate that this approach significantly enhances performance.
II Related Work
II-A Video frame interpolation
II-A1 General overview.
Numerous VFI solutions rely on optical flows to predict latent frames. Typically, these methods warp input frames forward or backward using flow calculated by off-the-shelf networks like  [14, 15, 16, 9] or self-contained flow estimators like [7, 17, 18]. Networks then refine the warped frame to improve visual quality. SuperSlomo [6] uses a linear combination of bi-directional flows for intermediate flow estimation and backward warping. DAIN [1] introduces a depth-aware flow projection layer for advanced intermediate flow estimation. AdaCoF [19] estimates kernel weights and offset vectors for each target pixel, while BMBC [20] and ABME [21] refine optical flow estimation. Large motion interpolation is addressed by XVFI [22] through a recursive multi-scale structure. VFIFormer [23] employs Transformers to model long-range pixel correlations. IFRNet [24], RIFE [7], and UPR-Net [25] employ efficient pyramid network designs for high-quality, real-time interpolation, with IFRNet and RIFE using leakage distillation losses for flow estimation. Recently, more advanced network modules and operations are proposed to push the upper limit of VFI performance, such as the transformer-based bilateral motion estimator of BiFormer [26], a unifying operation of EMA-VFI [17] to explicitly disentangle motion and appearance information, and bi-directional correlation volumes for all pairs of pixels of AMT [18]. On the other hand, SoftSplat [27] and M2M [28] actively explore the forward warping operation for VFI. More recently, diffusion-based video frame interpolators, such as LDMVFI [29] and SVDKFI [30], have leveraged generative priors to enhance visual quality but still face challenges such as high computational costs and slow runtimes.
@@ II-A Video frame interpolation
II-A2 Learning paradigms.
One major thread of VFI methods train networks on triplet of frames, always predicting the central frame. Iterative estimation is used for interpolation ratios higher than $\times$2. This fixed-time method often accumulates errors and struggles with interpolating at arbitrary continuous timesteps. Hence, models like SuperSloMo [6], DAIN [1], BMBC [20], EDSC [47], RIFE [7], IFRNet [24], EMA-VFI [17], and AMT [18] have adopted an arbitrary time interpolation paradigm. While theoretically superior, the arbitrary approach faces challenges of more complicated time-to-position mappings due to the velocity ambiguity, resulting in blurred results. This study addresses velocity ambiguity in arbitrary time interpolation and offers solutions.
Prior work by Zhou et al. [48] identified motion ambiguity and proposed a texture consistency loss to implicitly ensure interpolated content resemblance to given frames. In contrast, we explicitly address velocity ambiguity and propose solutions. These innovations not only enhance the performance of arbitrary time VFI models but also offer advanced manipulation capabilities. Additionally, we demonstrate that leveraging information from nearby frames through a multi-frame refiner module, combined with continuous indexing map estimation, can further improve interpolation quality.
II-A3 Segment anything
The emergence of Segment Anything Model (SAM) [11] has marked a significant advancement in the realm of zero-shot segmentation, enabling numerous downstream applications including video tracking and segmentation [49], breakthrough mask-free inpainting techniques [50], and interactive image description generation [51]. By specifying the distance indexing individually for each segment, this work introduces a pioneering application to this growing collection: Manipulated Interpolation of Anything.
III Velocity Ambiguity
@@
So far, we have been discussing the ambiguity for the fixed time interpolation paradigm, in which $t$ is set by default to $0.5$. For arbitrary time interpolation, the ambiguity becomes more pronounced: Instead of predicting a single timestep, the network is expected to predict a continuum of timesteps between 0 and 1, each having a multitude of possibilities. This further complicates learning. Moreover, this ambiguity is sometimes referred to as mode averaging, which has been studied in other domains [52]. See Appendix for details.
IV Disambiguation Strategies
@@ IV-A Distance indexing
During inference, the algorithm does not have access to the exact distance map since $I_{t}$ is unknown. In practice, we notice it is usually sufficient to provide a uniform map $D_{t}=t$, similar to time indexing. Physically this encourages the model to move each object at constant speeds along their trajectories. We observe that constant speed between frames is a valid approximation for many real-world situations. In Section VI, we show that even though this results in pixel-level misalignment with the ground-truth frames, it achieves significantly higher perceptual scores and is strongly preferred in the user study. Precise distance maps can be computed from multiple frames, which can potentially further boost the performance. See a detailed discussion in Appendix.
IV-B Iterative reference-based estimation
@@
Importantly, in every iteration, we consistently use the starting and ending frames as reliable appearance references, preventing divergence of uncertainty in each step. By dividing a long step into shorter steps, the uncertainty in each step is reduced, as shown in Fig. 3 (b). While fixed time models also employ an iterative method in a bisectioning way during inference, our strategy progresses from near to far, ensuring more deterministic trajectory interpolation. This reduces errors and uncertainties tied to a single, long-range prediction. See more on the rationale for solving ambiguities in Appendix.
Figure 5: Multi-frame fusion architecture with continuous map estimator.
V Leveraging Information from Nearby Frames
In this section, we first introduce a continuous indexing map estimator to achieve pixel-wise distance map estimation. Next, we demonstrate how to reuse the original interpolation architecture to fuse nearby frames for enhancing interpolation quality. The complete framework is illustrated in Fig. 5.
V-A Pixel-wise distance map estimation
To achieve a continuous indexing map estimator, we adopt the pretrained model introduced in CPFlow [12]. CPFlow takes a sequence of images as input (four in our case) and predicts the optical flow from the initial frame to any arbitrary timestamp within the sequence. Specifically, it models each pixel’s motion trajectory using cubic B-splines, enabling dense and temporally continuous flow estimation. Given normalized time $t\in[0,1]$, the displacement of a pixel is defined as:
${\mathbf{V}_{0\to t}=\sum_{i=0}^{N-1}B_{i,k}(t)\mathbf{P}_{i},}$ (12)
where $\mathbf{P}_{i}$ are learnable control points and $B_{i,k}(t)$ are spline basis functions defined recursively [12]. To enhance temporal consistency, the model employs a Neural Ordinary Differential Equation (NODE) module in combination with ConvGRU, where the hidden feature state $h(t)$ evolves according to:
[…]
The hidden state $h(t)$ is refined with frame-specific features $\varepsilon_{t}$ via ConvGRU: $\tilde{h}(t)=\text{ConvGRU}(h(t),\varepsilon_{t})$. The model then computes multi-scale correlation volumes between the reference feature $\tilde{h}(0)$ and $\tilde{h}(t)$, denoted by $C(t)=\text{Corr}(\tilde{h}(0),\tilde{h}(t))$. These correlations are used in an iterative decoder to update the control points of the spline as:
${\mathbf{P}_{i}^{(s+1)}=\mathbf{P}_{i}^{(s)}+\Delta\mathbf{P}_{i}^{(s)}(C(t)),}$ (14)
where $s$ is the iteration index. The final continuous optical flow $\mathbf{V}_{0\rightarrow t}$ is reconstructed from the refined control points using the spline formulation above. This continuous formulation allows the model to produce high-fidelity indexing maps using Eq. 7 for arbitrary interpolation times $D_{t}=\left\|\mathbf{V}_{0\to t}\right\|\cos{\theta}/\left\|\mathbf{V}_{0\to 1}\right\|$.
V-B Multi-frame refiner
We further design a multi-frame fusion module to leverage the relevant pixel appearance information beyond the two adjacent frames. As a brief review, flow-based VFI models [7] typically adhere to the following formulation:
${I_{t}^{+},I_{t}^{-},M=\mathcal{F}(I_{0},I_{1},D_{t}),}$ (15)
where $M$ is a one-channel blending mask, and $I_{t}^{+}$ and $I_{t}^{-}$ denote the warped images derived from $I_{0}$ and $I_{1}$ using the predicted optical flows. The final interpolated frame $I_{t}$ is obtained by:
${I_{t}=M\odot I_{t}^{+}+(1-M)\odot I_{t}^{-}}.$ (16)
An additional residual term is often included which is omitted here. To adapt the model to accept four consecutive frames as input while maintaining plug-and-play compatibility with various VFI models, we introduce a simple yet effective framework: we create a trainable copy of the two-frame VFI model $\mathcal{F^{\prime}}$, which accepts the additional frames $I_{-1}$ and $I_{2}$ along with a new distance map $D_{t}^{\prime}$ computed relative to $I_{-1}$ and $I_{2}$. Additionally, the initially interpolated frame $I_{t}$ is provided, enabling the network to refine the result by utilizing this supplementary information:
${I_{t}^{{}^{\prime}+},I_{t}^{{}^{\prime}-},M^{\prime}=\mathcal{F^{\prime}}(I_{-1},I_{2},D^{\prime}_{t},I_{t}),}$ (17)
where $I_{t}^{{}^{\prime}+},I_{t}^{{}^{\prime}-}$ are warped version of $I_{-1},I_{2}$ respectively. $M^{\prime}=[M_{1},M_{2},M_{3}]$ is a three-channel blending mask such that $M_{1}+M_{2}+M_{3}=1$ at each pixel, corresponding to $I_{t}^{{}^{\prime}+}$, $I_{t}$, and $I_{t}^{{}^{\prime}-}$. The final refined frame $I^{\prime}_{t}$ is then computed as:
${I^{\prime}_{t}=M_{1}\odot I_{t}^{{}^{\prime}+}+M_{2}\odot I_{t}+M_{3}\odot I_{t}^{{}^{\prime}-}}.$ (18)
Notice that we directly use the interpolated frame $I_{t}$ instead of latent features as network input to ensure compatibility with different VFI models.
@@ V-B Multi-frame refiner
Figure 6: Comparison of qualitative results. $[T]$: original arbitrary time VFI models using time indexing. $[D]_{u}$: models trained using our distance indexing, then inference using uniform maps. $[T,R]$: models using time indexing with iterative reference-based estimation. $[D,R]_{u}$: models trained using both distance indexing and iterative reference-based estimation, then inference using uniform maps.
V-C Model tuning
During training, we first freeze the parameters of the two-frame VFI model $\mathcal{F}$ and the continuous map estimator. Only the refiner $\mathcal{F^{\prime}}$ is updated during this stage, using the same training configuration as the original model. This setup enables the refiner to learn how to enhance interpolation by leveraging information from more distant adjacent frames. Next, we freeze the parameters of the pretrained map estimator and jointly optimize both the VFI model $\mathcal{F}$ and the proposed refiner $\mathcal{F^{\prime}}$. This approach enables the entire system to adapt to indexing maps derived from optical flow predicted by CPFlow. We experimented with various training strategies and loss functions to evaluate their effectiveness (see Section VI-H).
VI Experiment
VI-A Implementation
@@
We leveraged the plug-and-play nature of our distance indexing and iterative reference-based estimation strategies to seamlessly integrate them into influential arbitrary time VFI models such as RIFE [7] and IFRNet [24], and state-of-the-art models including AMT [18] and EMA-VFI [17]. We adhere to the original hyperparameters for each model for a fair comparison and implement them with PyTorch [53]. For training, we use the septuplet dataset from Vimeo90K [54]. The septuplet dataset comprises 91,701 seven-frame sequences at $448\times 256$, extracted from 39,000 video clips. For evaluation, we use both pixel-centric metrics like PSNR and SSIM [55], and perceptual metrics such as reference-based LPIPS [56] and non-reference NIQE [57]. Concerning the iterative reference-based estimation strategy, $D_{ref}$ during training is calculated from the optical flow derived from ground-truth data at a time point corresponding to a randomly selected reference frame, like $t/2$. In the inference phase, we similarly employ a uniform map for reference, for example, setting $D_{ref}=t/2$.
VI-B Qualitative comparison
@@
VI-C Quantitative comparison
VI-C1 Convergence curves
To further substantiate the efficacy of our proposed strategies, we also conducted a quantitative analysis. Fig. 7 shows the convergence curves for different model variants, i.e., $[T]$, $[D]$, and $[D,R]$. The observed trends are consistent with our theoretical analysis from Section IV, supporting the premise that by addressing velocity ambiguity, both distance indexing and iterative reference-based estimation can enhance convergence limits.
VI-C2 Comparison on Vimeo90K septuplet dataset
In Table I, we provide a performance breakdown for each model variant. The models $[D]$ and $[D,R]$ in the upper half utilize ground-truth distance guidance, which is not available at inference in practice. The goal here is just to show the achievable upper-bound performance. On both pixel-centric metrics such as PSNR and SSIM, and perceptual measures like LPIPS and NIQE, the improved versions $[D]$ and $[D,R]$ outperform the base model $[T]$. Notably, the combined model $[D,R]$ using both distance indexing and iterative reference-based estimation strategies performs superior in perceptual metrics, particularly NIQE. The superior pixel-centric scores of model $[D]$ compared to model $[D,R]$ can be attributed to the indirect estimation (2 iterations) in the latter, causing slight misalignment with the ground-truth, albeit with enhanced details.
TABLE I: Comparison on Vimeo90K septuplet dataset. $[T]$ denotes the method trained with traditional arbitrary time indexing paradigm. $[D]$ and $[R]$ denote the distance indexing paradigm and iterative reference-based estimation strategy, respectively. $[R]$ uses 2 iterations by default. $[\cdot]_{u}$ denotes inference with uniform map as time indexes. We utilize the first and last frames as inputs to predict the rest five frames. The
bold font
denotes the best performance in cases where comparison is possible. While the gray font indicates that the scores for pixel-centric metrics, PSNR and SSIM, are not calculated using strictly aligned ground-truth and predicted frames.